BaconUp:

Solve the equation. Then check your solution. (-2 1/7)p=-3 a. 1 2/5 b. 6 3/7 c. -1 2/5 d. -6 3/7

BaconUp:

please help

tigerlover:

p in (-oo:+oo) ((-2*1)/7)*p = -3 // + 3 ((-2*1)/7)*p+3 = 0 (-2/7)*p+3 = 0 3-2/7*p = 0 // - 3 -2/7*p = -3 // : -2/7 p = -3/(-2/7) p = 21/2 p = 21/2

Vocaloid:

@BaconUp start by converting -2 and 1/7 to a mixed number then divide both sides by that mixed number to isolate p